# Towards Operational Automated Greenhouse Gas Plume Detection

### v1

@@ 7 Discussion
Background Sampling: Background samples play a crucial role in constructing and validating a robust plume detector. Collecting a representative negative class dataset (i.e. the background) is as important as collecting a representative positive class dataset (i.e. plumes), and failure to do so will result in poor generalization. For example, sampling an equal number of plume and background tiles for model development results in a biased representation of the true observed distribution, where plume pixels are a distinct minority (Růžička et al., 2023). Sampling such few background tiles also misses many of the rare false enhancements that drive most false positive detections. Therefore, a model trained and evaluated on datasets with equal quantities of the plume and background will appear to perform well, but it will generalize poorly operationally when encountering the true observed distribution and the most challenging false enhancements. We illustrate this in detail in Suppl. Section LABEL:apx:bias_sample and Figure LABEL:fig:instance_tilebias.
@@
Our work addresses this potential issue by sampling as many background tiles from scenes as possible. As described in Section 3 and Table 5, the plume-to-background ratio of the AVIRIS-NG dataset used to train our models was less than 1:13, an unbalanced sampling that more accurately reflects the observed rarity of high concentrations, which allows our models to perform well when evaluated on entire scenes without tile sampling in Section 5.1.
Tilewise and Scenewise Capabilities: Even with proper sampling, tilewise plume detection performance metrics (Figure 8) tend to be optimistic estimates of scenewise plume detection accuracy (Figure 6). This gap between tilewise and scenewise performance is exacerbated by any undersampling of the background class as discussed previously. Since operational plume detection systems are deployed on entire scenes and not sampled tiles, it is crucial to report scenewise capabilities (as in Section 5.1) in addition to any tilewise capabilities, which remain useful for comparing model performance (as in Section 5.2) but not for practical evaluation of operational applicability.
Unambiguous Detection and Segmentation Metrics: While most prior work reports the metric for which a model was optimized, measuring the operational performance of a proposed plume detector requires measuring both instance detection and pixelwise segmentation performance as defined in Section 2.5. As Figure 6 illustrates, a model which performs poorly in pixelwise plume segmentation metrics (e.g., our tilewise-trained model) can significantly outperform comparable models in plume instance detection metrics. Recognizing that certain use cases may prioritize either segmentation (e.g. plume masking for emission rate estimation) or detection (e.g. identifying sources for review), reporting both metrics is necessary for complete operational evaluation.
Capturing detection and segmentation accuracy with respect to false positive and false negative prediction errors is also important to provide an unbiased view of a proposed model’s capabilities. “Coupled” metrics such as traditional classification accuracy can provide a misleading view of performance, especially when samples are imbalanced. In cases where a single metric is necessary for comparison, F- scores, where weighs recall relative to precision, may be appropriate in addition to precision and recall.
Replicable Data Collection & Sampling: Assessing the operational effectiveness of a ML-driven GHG plume detection approach requires detailed accounts that describe the assumptions and practices involved in data collection, tile sampling and cross-validation. Specifics describing how samples were collected and the scope and diversity of both the plume class and the background class are necessary to determine if they are sufficiently representative to construct the model and assess its effectiveness and the extent which known sources of bias might impact reproducibility on similar data, and potential replicability in broader operational settings. For instance, if a proposed approach does not specify whether sampled tiles are allowed to overlap within scenes, across distinct scenes, or between plume and background samples, we cannot assess the impact of spatiotemporal bias on the associated model’s predictions, nor can we replicate results on similar data. To-date, detailed data collection and model validation procedures have been sorely lacking in the literature; we aimed to address this by thoroughly describing data sampling strategies in Section 3, supported by Suppl. Sections LABEL:apx:datainfo and LABEL:apx:qc.
Observational Validation of WRF-LES-based Plume Detectors: A common theme among proposed plume detection approaches is to construct synthetic GHG plume observations by injecting plumes with known characteristics generated by the WRF-LES system into plume-free “background/noise” retrieval products derived from data captured by the target platform. However, all such works in the literature are limited in evaluating their WRF-LES-based detection systems on observed data captured by the target remote sensing platforms for operational deployment. Prior works that only evaluate their WRF-LES-trained models on other WRF-LES plumes (Jongaramrungruang et al., 2022; Bruno et al., 2023) are difficult to discern for operational feasibility due to the significant differences between plumes generated by WRF-LES and real-world plumes. Other prior works do discuss results of their WRF-LES-trained models on observed methane plumes (Joyce et al., 2022; Radman et al., 2023; Xu et al., 2024); however, they only report a couple dozen true positive detections given small scenes containing previously identified plumes, or perform emission rate quantification on manually detected plumes. The omission of false positive or false negative detections on a large number of instrument-observed scenes makes it difficult to determine the capabilities of these methods in an operational setting.
In order to determine the capabilities of WRF-LES trained models when applied to observational data, we reproduced the plume detection results of Jongaramrungruang et al. (2022) and compared the capabilities of the WRF-LES-trained models to AVIRIS-NG-trained models in Suppl. Section LABEL:apx:bias_distrib. The results demonstrate poor generalization, with the WRF-LES-trained models failing to perform beyond a 0.3 F1-Score in tilewise instance detection on AVIRIS-NG test datasets despite reporting a nearly 1.0 F1-Score on the WRF-LES test dataset. This demonstrates that discrepancies between WRF-LES simulations and CMF GHG retrievals pose a significant challenge to the operational deployment of models trained solely on WRF-LES plumes, despite being injected on real-world backgrounds. Such discrepancies include the lack of changes in wind speed and direction at different times and altitudes, as well as the inability to simulate non-point-source diffuse emissions, which make up a significant portion of real world emissions and the AVIRIS-NG dataset.
Avoiding Spatiotemporal Bias and Data Leakage: CNN-based plume detectors are prone to memorizing, or “overfitting” to, site-specific features that occur in spatially overlapping observations of the same site, or from overlapping image tiles extracted from the same retrieval image scene. When overlapping images are present in both training and test sets, data leakage occurs. Validation procedures corrupted by data leakage generally produce highly optimistic validation metrics, as the model is able to “cheat” on test samples that overlap training samples with spatially correlated features. For example, Schuit et al. (2023) reports precision and recall scores over 0.97 using a CNN to distinguish tiles of labeled plume and non-plume tiles from TROPOMI XCH4 products between 2018-2020. However, the plume samples are derived from observations of only 60 regions with persistent emission sources, making it likely that many of the plumes represent repeat observations of the same emission sources. When tiles with 50% overlap are sampled from each TROPOMI image, then randomly split into training and test sets without stratification, data leakage inevitably occurs, likely accounting for the abnormally high precision and recall detection performances. When applied to TROPOMI XCH4 products from 2021, this work reports a more realistic precision score of 0.61 for confident plume detections and 0.76 for both confident and ambiguous detections. However, these results are reported after an undefined filtering process and calculated from a post hoc analysis of detection results instead of a previously labeled dataset, therefore omitting false negative analysis.
@@
Our work makes a significant effort to control for spatiotemporal bias and prevent data leakage in our tile sampling and cross validation procedures (described in Section 3) by ensuring that spatially overlapping observations are assigned to either the training or the test dataset, but not both. As a result, a plume nor its source location (whether emitting or not) will never appear in both the training and test datasets. We also demonstrate the potential impacts of data leakage in Suppl. Section LABEL:apx:bias_stratify by inducing data leakage with a random non-stratified train-test split and showing its results. One such experiment with data from the Four Corners campaign showed that failing to prevent data leakage resulted in a misleading inflation of the model F1-Score by 83%.
@@
Spatiotemporal bias and data leakage also likely contribute to the high overall detection accuracies reported in WRF-LES-based results, as most prior works inject multiple versions of WRF-LES generated plumes into the same observed “background” retrieval image, or worse, repeatedly inject a single WRF-LES plume instance into multiple distinct backgrounds. Allowing these samples to cross training and test boundaries will introduce severe validation bias, as the model will be tested on plumes and backgrounds identical to those in the training set. The former is especially common in prior works, with many training datasets composed of many WRF-LES simulated plumes injected into a smaller collection of background images (Joyce et al., 2022; Radman et al., 2023; Bruno et al., 2023; Jongaramrungruang et al., 2022). If the same plume-background images appear in both the training and test datasets, it increases the likelihood that the CNN learns to make plume detections based on background surface features seen during training (e.g. roads, infrastructure, geographic features) instead of the plume itself (by “overfitting”). Deliberate care must be taken to ensure background tiles do not leak between the training and test datasets.
@@ 8 Conclusions
While previously published work claimed promising proof-of-concept results using ML-driven approaches for plume detection, detailed review revealed a variety of methodological issues that lead to significant biases and overly optimistic results for real-world deployments. Furthermore, due to the absence of crucial details regarding data collection, model optimization, and/or validation procedures, the majority of published results are not reproducible experimentally. Prior efforts all lack rigorous validation procedures to ensure that trained models generalize on independent and identically distributed observational data that generally lead to optimistic results driven by data and model-driven biases. Consequently, even in cases where the published results are feasible to reproduce, successfully replicating published accuracies on distinct but analogous data is unlikely. Here, we take steps to address these gaps in the literature, with the goal of pushing towards operational deployments of automated plume detection models.
@@
AVIRIS-NG and EMIT datasets as described, both for tilewise model training and scenewise evaluation, are available at https://ml.jpl.nasa.gov/datasets/opsghgml/. GAO datasets are not available for public access due to data controls, and only evaluation results are presented as reference. Dataset sampling and preprocessing code, as well as model training, inference, and evaluation code, is open source and available at https://github.com/JPLMLIA/operational-ghg.

### v2

@@ 6 Discussion
6.1 Addressing Common Limitations
6.1.1 Biases in Data Sampling
@@
Background Sampling: Background samples play a crucial role in constructing and validating a robust plume detector. Collecting a representative negative class dataset (i.e. the background) is as important as collecting a representative positive class dataset (i.e. plumes), and failure to do so will result in poor generalization. Our work addresses this potential issue by sampling as many background tiles from scenes as possible. As described in Section 3 and Table 1, the plume-to-background ratio of the AVIRIS-NG dataset used to train our models was less than 1:13, an unbalanced sampling that more accurately reflects the observed rarity of high concentrations, which allows our models to perform well when evaluated on entire scenes without tile sampling in Section 5.1.
By contrast, sampling an equal number of plume and background tiles for model development, as in Růžička et al. (2023), results in a biased representation of the true observed distribution, in which plume pixels are a distinct minority. Sampling such few background tiles also misses many of the rare false enhancements that drive most false positive detections. Therefore, a model trained and evaluated on datasets with equal quantities of the plume and background will appear to perform well, but it will generalize poorly operationally when encountering the true observed distribution and the most challenging false enhancements. We illustrate this in detail in Suppl. Section LABEL:apx:bias_sample and Figure LABEL:fig:instance_tilebias.
@@
Avoiding Spatiotemporal Bias and Data Leakage: CNN-based plume detectors are prone to memorizing, or “overfitting” to, site-specific features that occur in spatially overlapping observations of the same site, or from overlapping image tiles extracted from the same retrieval image scene. When overlapping images are present in both training and test sets, data leakage occurs. Validation procedures corrupted by data leakage generally produce highly optimistic validation metrics, as the model is able to “cheat” on test samples that overlap training samples with spatially correlated features. Our work makes a significant effort to control for spatiotemporal bias and prevent data leakage in our tile sampling and cross validation procedures (described in Section 3) by ensuring that spatially overlapping observations are assigned to either the training or the test dataset, but not both. As a result, a plume nor its source location (whether emitting or not) will never appear in both the training and test datasets.
By contrast, Schuit et al. (2023) reports precision and recall scores over 0.97 using a CNN to distinguish tiles of labeled plume and non-plume tiles from TROPOMI XCH4 products between 2018-2020. However, the plume samples are derived from observations of only 60 regions with persistent emission sources, making it likely that many of the plumes represent repeat observations of the same emission sources. When tiles with 50% overlap are sampled from each TROPOMI image, then randomly split into training and test sets without stratification, data leakage inevitably occurs, likely accounting for the abnormally high precision and recall detection performances. Schuit et al. (2023) reports a more realistic precision score of 0.61 for confident plume detections and 0.76 for both confident and ambiguous detections on TROPOMI XCH4 products from 2021. However, these results are reported after an undefined filtering process and calculated from a post hoc analysis of detection results instead of a previously labeled dataset, therefore omitting false negative analysis.
We further demonstrate the potential impacts of data leakage in Suppl. Section LABEL:apx:bias_stratify by inducing data leakage with a random non-stratified train-test split and showing its results. One such experiment with data from the Four Corners campaign showed that a random train-test split resulted in an inflated F1 Score of 0.75, whereas the stratified train-test split where data leakage was mitigated resulted in the actual F1 Score of 0.41. Other experiments with the CalCH4 (random 0.92; stratified 0.80) and Permian (random 0.86; stratified 0.81) datasets showed similar results, whereas the COVID (random 0.79; stratified 0.81) did not show this effect.
Replicable Data Collection & Sampling: Assessing the operational effectiveness of a ML-driven GHG plume detection approach requires detailed accounts that describe the assumptions and practices involved in data collection, tile sampling and cross-validation. Specifics describing how samples were collected and the scope and diversity of both the plume class and the background class are necessary to determine if they are sufficiently representative to construct the model and assess its effectiveness and the extent which known sources of bias might impact reproducibility on similar data, and potential replicability in broader operational settings. For instance, if a proposed approach does not specify whether sampled tiles are allowed to overlap within scenes, across distinct scenes, or between plume and background samples, we cannot assess the impact of spatiotemporal bias on the associated model’s predictions, nor can we replicate results on similar data. To-date, detailed data collection and model validation procedures have been sorely lacking in the literature; we aimed to address this by thoroughly describing data sampling strategies in Section 3, supported by Suppl. Sections LABEL:apx:datainfo and LABEL:apx:qc. We also release our dataset of sampled tiles for AVIRIS-NG and EMIT, as described in the Data and Code Availability statement.
6.1.2 Informative Performance Metrics
Even with proper sampling, tilewise plume detection performance metrics (Figure 7) tend to be optimistic estimates of scenewise plume detection accuracy (Figure 5). This gap between tilewise and scenewise performance is exacerbated by any undersampling of the background class, as discussed previously. Since operational plume detection systems are deployed on entire scenes and not sampled tiles, it is crucial to report scenewise performance metrics (as in Section 5.1) in addition to any tilewise performance metrics. While tilewise metrics remain useful for comparing model performance (as in Section 5.3), they are insufficient for the practical evaluation of operational applicability.
Measuring the operational performance of a proposed plume detector requires measuring both instance detection and pixelwise segmentation performance as defined in Section 2.5. As Figure 5 illustrates, a model which performs poorly in pixelwise plume segmentation metrics (e.g., our tilewise-trained model) can significantly outperform comparable models in plume instance detection metrics. Recognizing that certain use cases may prioritize either segmentation (e.g. plume masking for emission rate estimation) or detection (e.g. identifying sources for review), reporting both metrics is necessary for complete operational evaluation. This comprehensive approach addresses a common limitation in the literature, where most prior work only report the tilewise metrics for which a model was optimized.
Finally, capturing detection and segmentation accuracy with respect to false positive and false negative prediction errors is also important to provide an unbiased view of a proposed model’s capabilities. “Coupled” metrics such as traditional classification accuracy can provide a misleading view of performance, especially when samples are imbalanced. In cases where a single metric is necessary for comparison, F- scores, where weighs recall relative to precision, may be appropriate in addition to precision and recall.
6.1.3 The Simulation Gap of WRF-LES Datasets
A common theme among proposed plume detection approaches is to construct synthetic GHG plume observations by injecting plumes with known characteristics generated by the WRF-LES system into plume-free “background/noise” retrieval products derived from data captured by the target platform. This approach may be attractive because “perfect” label masks can be generated for these simulated plume based on known concentrations. However, all such works in the literature are limited in evaluating their WRF-LES-based detection systems on observed data captured by the target remote sensing platforms for operational deployment.
Prior works that only evaluate their WRF-LES-trained models on other WRF-LES plumes (Jongaramrungruang et al., 2022; Bruno et al., 2024) are difficult to discern for operational feasibility due to the significant differences between plumes generated by WRF-LES and real-world plumes. Other prior works do discuss results of their WRF-LES-trained models on observed methane plumes (Joyce et al., 2022; Radman et al., 2023; Xu et al., 2024); however, they only report a couple dozen true positive detections given small scenes containing previously identified plumes, or perform emission rate quantification on manually detected plumes. The omission of false positive or false negative detections on a large number of instrument-observed scenes makes it difficult to determine the capabilities of these methods in an operational setting.
In order to determine the capabilities of WRF-LES trained models when applied to observational data, we reproduced the plume detection results of Jongaramrungruang et al. (2022) and compared the capabilities of the WRF-LES-trained models to AVIRIS-NG-trained models in Suppl. Section LABEL:apx:bias_distrib. The results demonstrate poor generalization, with the WRF-LES-trained models failing to perform beyond a 0.3 F1-Score in tilewise instance detection on AVIRIS-NG test datasets despite reporting a nearly 1.0 F1-Score on the WRF-LES test dataset. This demonstrates that discrepancies between WRF-LES simulations and CMF GHG retrievals pose a significant challenge to the operational deployment of models trained solely on WRF-LES plumes, despite being injected on real-world backgrounds. Discrepancies in this case include the lack of changes in wind speed and direction over time and as the plume rises in altitude, as well as the lack of diffuse emission simulations, which make up a significant portion of real world emissions and the AVIRIS-NG dataset.
@@ 6.1.3 The Simulation Gap of WRF-LES Datasets
Spatiotemporal bias and data leakage also likely contribute to the high overall detection accuracies reported in WRF-LES-based results, as most prior works inject multiple versions of WRF-LES generated plumes into the same observed “background” retrieval image, or worse, repeatedly inject a single WRF-LES plume instance into multiple distinct backgrounds. Allowing these samples to cross training and test boundaries will introduce severe validation bias, as the model will be tested on plumes and backgrounds identical to those in the training set. The former is especially common in prior works, with many training datasets composed of many WRF-LES simulated plumes injected into a smaller collection of background images (Joyce et al., 2022; Radman et al., 2023; Bruno et al., 2024; Jongaramrungruang et al., 2022). If the same plume-background images appear in both the training and test datasets, it increases the likelihood that the CNN learns to make plume detections based on background surface features seen during training (e.g. roads, infrastructure, geographic features) instead of the plume itself (by “overfitting”). Deliberate care must be taken to ensure background tiles do not leak between the training and test datasets.
@@ 7 Conclusions
While previously published work claimed promising proof-of-concept results using ML-driven approaches for plume detection, our review revealed that a variety of methodological issues discussed above, including undersampling the background, the simulation gap between WRF-LES and observed plumes, and spatiotemporal biases, likely explain overly optimistic performance metrics compared to real-world deployments. Furthermore, due to the absence of crucial details regarding data collection, model optimization, and/or validation procedures, the majority of published results are not reproducible experimentally. Prior efforts all lack rigorous validation procedures to ensure that trained models generalize on independent and identically distributed observational data that generally lead to optimistic results driven by data and model-driven biases. Consequently, even in cases where the published results are feasible to reproduce, successfully replicating published accuracies on distinct but analogous data is unlikely. Here, we take steps to address these gaps in the literature, with the goal of pushing towards operational deployments of automated plume detection models.
@@
AVIRIS-NG and EMIT datasets as described are available at https://doi.org/10.5281/zenodo.19011045. GAO datasets are not available for public access due to data controls, and only evaluation results are presented as reference. Dataset sampling and preprocessing code, as well as model training, inference, and evaluation code, is open source and available at https://github.com/JPLMLIA/operational-ghg.